# TelePhantom: A User-Friendly Teleoperation System with Virtual Assistance for Enhanced Effectiveness

### v1

@@ Abstract
Dexterous manipulation is a critical area of robotics. In this field, teleoperation faces three key challenges: user-friendliness for novices, safety assurance, and transferability across different platforms. While collecting real robot dexterous manipulation data by teleoperation to train robots has shown impressive results on diverse tasks, due to the morphological differences between human and robot hands, it is not only hard for new users to understand the action mapping but also raises potential safety concerns during operation. To address these limitations, we introduce TelePhantom. This teleoperation system offers real-time visual feedback on robot actions based on human user inputs, with a total hardware cost of less than \$1,000. TelePhantom allows the user to see a virtual robot that represents the outcome of the user’s next movement. By enabling flexible switching between command visualization and actual execution, this system helps new users learn how to demonstrate quickly and safely. We demonstrate its superiority over other teleoperation systems across five tasks, emphasize its ease of use, and highlight its ease of deployment across diverse input sensors and robotic platforms. We will release our code and a deployment document on our website https://telephantom.github.io/.
@@
Figure 1: TelePhantom is a user-friendly teleoperation system enabling the real-time virtual preview before robot execution.
@@ I Introduction
Figure 2: Overview of the TelePhantom System Architecture: Our system consists of three main components: (1) Various input devices including RGBD cameras, IMU mocap suits, VR headsets, mocap gloves, AR hand tracking, and EMG armbands for capturing human motion; (2) A central processing pipeline that uses SMPL-X as a unified interface to process wrist and hand poses data, feeding into our Phantom System with joint-to-joint mapping and non-collision retargeting network; (3) Support for different robot platforms as output devices for executing the mapped movements.
@@
Teleoperation has long been recognized as an essential component in gathering on-robot data for learning from demonstrations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. An intuitive and user-friendly teleoperation system is crucial for collecting high-quality, diverse, and scalable data. However, effectively controlling robots with dexterous hands remains a significant challenge. The fundamental differences between human and robotic hands make the direct mapping of human movements to robotic actions exceptionally complex. These inherent structural and functional differences, combined with the precision required for precise manipulation tasks, often result in inefficient and potentially compromised data collection processes.
@@
Existing teleoperation systems face three significant challenges. First, novice users often struggle to control the robot effectively due to the lack of intuitive feedback on how their commands translate to robot actions, especially during complex manipulation tasks. Second, without proper safeguards and real-time guidance, users may inadvertently issue unsafe commands that could damage the robot or its surroundings. Third, most current systems are tightly coupled with specific input devices or robot platforms, limiting their adaptability and broader application. These challenges significantly impact the quality and efficiency of data collection for robot learning.
To address these challenges, we introduce TelePhantom, a teleoperation system that delivers precise control while maintaining affordability through its sub-\$1,000 hardware implementation. At its core, TelePhantom provides an interface where a virtual robotic arm spatially overlaps with the physical arm, allowing users to preview and confirm intended movements through an external input device (e.g., a foot pedal) before execution on the physical system.
Building upon this core design, our modular framework (Figure 2) addresses key challenges in human-robot motion mapping through the seamless integration of multiple components. Our system architecture utilizes IMU-based motion capture systems and mocap gloves to estimate wrist (Section III-A) and hand poses (Section III-B1). We use the SMPL-X standard (Section III-A1) to bridge all modules together, enabling flexible integration while maintaining system cohesion. These inputs are mapped to the robotic hand using our joint-to-joint retargeting approach (Section III-B2) and non-collision retargeting module (Section III-C). We design each module to be independent yet interoperable, enabling flexible system configuration while maintaining robust performance.
We present the major contributions of TelePhantom to advance the field of teleoperation:
  • Low-Cost System: We introduce an economical yet high-performance teleoperation system that integrates inertial motion capture technology and mocap gloves, with a total hardware cost under \$1,000, making it affordable for research.
@@
  • Interactive Visual Assistance: We develop a visualization interface that enables users to preview intended actions before execution, enhancing teleoperation precision and accessibility for both novice and experienced users during complex tasks.
  • Versatile Modules: We design a modular assisted teleoperation pipeline that bridges various input devices with different robot end-effectors through joint-to-joint mapping and non-collision retargeting. This cross-platform design enables the versatile module to be integrated with diverse input methods and robots, making it a flexible solution for various manipulation scenarios.
@@ II-A Robot Teleoperation Frameworks for Manipulation
The growing interest in training robots via imitation learning has driven the need for extensive, high-quality real robot datasets. Teleoperation provides an efficient method for demonstrating and recording intricate robotic tasks. Researchers have explored various teleoperation systems, including VR controllers [1, 2, 3], motion capture [12, 13, 14], wearable gloves [15, 11, 16, 17], exoskeletons [18, 19], and other low-cost devices, each offering distinct benefits in terms of accessibility, precision, and generalizability.
While vision-based teleoperation methods offer a relatively low-cost solution [6, 2, 3, 1], they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspace, and high computational demands. Although specialized hardware [4] can improve accuracy, it substantially increases system costs. Recent unified teleoperation frameworks attempt to address these limitations but often introduce new trade-offs. Many systems [3, 6] prioritize implementation simplicity by tracking only fingertip positions rather than enabling detailed joint-level control. Furthermore, their customization options typically require time-consuming manual parameter tuning that demands significant expertise.
Figure 3: Pipeline of Our Assisted Module: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides visual preview before physical execution. We achieve precise alignment between virtual and physical robots through AprilTag calibration.
@@ II-B Visual Feedback in Teleoperation Systems
Visual feedback plays a crucial role in teleoperation systems by providing users with spatial awareness and task-relevant information. Traditional approaches primarily rely on real-time camera feeds [20, 21], which can be limited by occlusions and restricted viewing angles. Moreover, the lack of preview capabilities for intended commands leaves users (especially novice users) uncertain about the robot’s response, which can increase error rates and user fatigue.
@@
To address these limitations, researchers have explored enhanced visualization techniques [22, 23, 24] that overlay virtual information onto the workspace view. While these approaches improve user perception by displaying planned trajectories and environmental constraints, they primarily focus on visualizing task-specific elements. This focus on task-specific visualization leaves users uncertain about how the robot will execute its commands.
III Modular Teleoperation Framework
Our teleoperation system consists of three main components: wrist pose estimation via IMU-based tracking, hand pose estimation using a motion capture glove, and non-collision retargeting for safe robot command generation. The following subsections detail each component of this pipeline.
III-A Wrist Pose Estimation
III-A1 SMPL-X Standard
@@
We adopt the SMPL-X standard [25] as our unified human representation. The SMPL-X model represents a kinematic tree with standardized joint coordinate systems, where each joint’s pose describes a rotation and translation relative to its parent frame. This hierarchical structure enables consistent pose representation across different motion capture devices and simplifies the integration with our TelePhantom system.
III-A2 IMU-based Tracking System
While many researchers adopt vision-based motion capture systems in teleoperation, these systems require multiple cameras, precise calibration, and controlled lighting conditions, which limits their practical deployment. IMU-based motion capture offers robust tracking regardless of visual conditions and occlusions.
@@ III-A2 IMU-based Tracking System
Our system uses IMU sensors located on the user’s palm, forearm, and upper arm. Given the three IMUs, we utilize the REBOCAP framework [26] to obtain the specific SMPL-X model of the human body [25, 27]. Following the SMPL-X convention, we define the world frame system at the midpoint between the feet. All positions and orientations in this paper are expressed in this coordinate system unless otherwise specified. After obtaining the SMPL-X model, we compute the wrist position $\mathbf{p}_{w}\in\mathbb{R}^{3}$ and orientation $\mathbf{R}_{w}\in SO(3)$ relative to this world frame.
@@
III-B Hand Pose Estimation
III-B1 Human Hand Joint Value Acquisition
Vision-based methods typically use deep learning models to predict 3D joint positions $\{p_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ from RGB or RGB-D images. Due to the projective nature of imaging, vision-based systems struggle with self-occlusion, forcing researchers to infer joint positions from priors rather than direct observation. In contrast, mocap gloves [28] offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors despite their higher hardware cost. For our teleoperation scenario, we opt for the mocap-based approach to ensure robust hand tracking.
III-B2 Joint-to-Joint Mapping
@@
III-C Non-Collision Retargeting Network
In robotic manipulation systems, ensuring non-collision during operation is crucial for both physical simulator accuracy and hardware safety. Self-collisions can lead to unstable behavior in simulators and, more critically, cause hardware damage in real-world deployments. This challenge is particularly significant for high-DoF systems like the LEAP hand [29], where collisions often occur due to differences in action spaces and imprecise retargeting. As shown in Fig.4, direct mapping from human hand postures to robotic hands using existing retargeting approaches like [6] can result in self-collisions, as demonstrated in Fig.4 (a). While traditional collision avoidance methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations.
To address these challenges, we develop a learning-based framework with two key components: (1) a Self-Collision Prediction Network (CPN) that predicts collision probabilities for each joint configuration and (2) a Configuration Correction Network (CCN) that maps invalid configurations to their closest valid counterparts. The CCN is trained with a composite loss function that balances configuration similarity and collision avoidance, ensuring both motion continuity and physical feasibility. As demonstrated in Fig. 4 (b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz). We provide the detailed technical implementation of this Pos-to-Pos Non-Collision Module in the Appendix.
(a) Current retargeting method [6] causing collision.
@@ III-C Non-Collision Retargeting Network
IV Phantom System
The phantom system aims to provide safe and intuitive robot teleoperation through a phantom-based motion preview mechanism. The process consists of three key steps: (1) the user’s motion commands are first rendered on a phantom robot, (2) the phantom is precisely aligned with the real robot using AprilTag markers [30], and (3) upon user confirmation via a foot pedal, the validated motion is executed by the physical robot. This preview-then-execute workflow, as shown in Fig. 3, enables users to verify and adjust planned motions before actual execution, significantly improving teleoperation safety and efficiency.
IV-A Why Phantom System Assistance?
IV-A1 Current Challenges
Traditional teleoperation approaches face significant challenges in data quality when collecting demonstrations for imitation learning. During complex manipulation tasks, users frequently make exploratory movements that contaminate the demonstration data with sub-optimal trajectories.
@@ IV-A1 Current Challenges
In our hanging insertion experiments, we observed that users using direct teleoperation required an average of 4 corrective movements per attempt. These adjustments typically involved probing movements to locate the socket edge and multiple approach angle refinements before successful insertion. Such trial-and-error behavior introduced significant noise into the collected datasets. This issue is particularly problematic for training imitation learning models, which benefit most from clean, purposeful demonstrations rather than trajectories cluttered with exploratory movements.
IV-A2 Proposed Solution
Our phantom-assisted teleoperation system addresses these data quality challenges through a novel preview mechanism. By visualizing a virtual “phantom” representation of the robot’s intended motion before executing physical movements, users can preview and refine their commands before committing to actual robot motion. This preview capability eliminates the need for exploratory movements in the physical space, as the real robot remains stationary until the user confirms the phantom pose. This design ensures that recorded trajectories contain only intentional, task-relevant actions without the noise of exploratory movements, directly benefiting downstream imitation learning applications with higher-quality demonstration data.
IV-B Technical Implementation of Phantom System
IV-B1 Sim-to-Real Alignment Using AprilTag
To maintain precise alignment between the phantom and physical robot across multiple camera viewpoints, we implement a robust calibration system using AprilTags [30]. Our approach establishes reliable reference frames through hand-eye calibration, and AprilTag pose estimation, linking both fixed and floating camera coordinate systems to the robot’s base frame.

### v4

# TelePreview: A User-Friendly Teleoperation System with Virtual Arm Assistance for Enhanced Effectiveness

@@ Abstract
Teleoperation provides an effective way to collect robot data, which is crucial for learning from demonstrations. In this field, teleoperation faces several key challenges: user-friendliness for new users, safety assurance, and transferability across different platforms. While collecting real robot dexterous manipulation data by teleoperation to train robots has shown impressive results on diverse tasks, due to the morphological differences between human and robot hands, it is not only hard for new users to understand the action mapping but also raises potential safety concerns during operation. To address these limitations, we introduce TelePreview. This teleoperation system offers real-time visual feedback on robot actions based on human user inputs, with a total hardware cost of less than \$1,000. TelePreview allows the user to see a virtual robot that represents the outcome of the user’s next movement. By enabling flexible switching between command visualization and actual execution, this system helps new users learn how to demonstrate quickly and safely. We demonstrate that it outperforms other teleoperation systems across five tasks, emphasize its ease of use, and highlight its straightforward deployment across diverse robotic platforms. We release our code and a deployment document on our website https://nus-lins-lab.github.io/telepreview/.
@@
Figure 1: TelePreview is a user-friendly teleoperation system enabling the real-time virtual preview before robot execution.
@@ I Introduction
Figure 2: Overview of the TelePreview System Architecture: Our system consists of three main components: (1) Various input devices, including RGB-D cameras, IMU mocap suits, VR headsets, mocap gloves, AR hand tracking, and EMG armbands for capturing human motion; (2) A central processing pipeline that uses SMPL-X as a standard to process wrist and hand pose data, feeding into our Preview System with joint-to-joint mapping and a non-collision retargeting network; (3) Support for different robot platforms as output devices for executing the mapped movements.
@@
Researchers have long recognized teleoperation as an essential component for gathering on-robot data used in learning from demonstrations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. An intuitive and user-friendly teleoperation system is crucial for collecting high-quality, diverse, and scalable data. However, effectively controlling robots with dexterous hands remains a significant challenge. The fundamental differences between human and robotic hands make the direct mapping of human movements to robotic actions exceptionally complex. These inherent structural and functional disparities, combined with the precision required for fine-grained tasks, often result in inefficient and potentially compromised data collection processes.
@@
Existing teleoperation systems face three significant challenges. First, new users often struggle to control the robot effectively due to the lack of intuitive feedback on how their commands translate to robot actions, especially during complex manipulation tasks. Second, without proper safeguards and real-time guidance, users may inadvertently issue unsafe commands that could damage the robot or its surroundings. Third, most current systems are tightly coupled with specific input devices or robot platforms, limiting their adaptability. These challenges significantly impact the quality and efficiency of data collection for robot learning.
To address these issues, we introduce TelePreview, a teleoperation system that delivers precise control while maintaining affordability through its sub-\$1,000 hardware setup, as shown in Fig. 2. At its core, TelePreview provides an interface that overlays a virtual robotic arm onto the real-world scene, enabling users to preview motions. Precisely, a foot pedal or other external signal IO allows for mode switching between preview (virtual-only) and align (physical action alignment) modes. In preview mode, users can verify and refine their intended actions with a virtual robot arm that is spatially aligned with the physical robot, ensuring that the final commands are safe and accurate once applied in the real world.
We present the significant contributions of TelePreview to advance teleoperation:
@@
  • Interactive Visual Assistance: We develop a visualization interface that enables users to preview intended actions before execution, enhancing teleoperation precision and accessibility for both new and experienced users during complex tasks.
  • Low-Cost System: We introduce an economical yet high-performance teleoperation solution that integrates inertial motion capture technology and mocap gloves at a total hardware cost of under \$1,000, making it affordable for research.
  • Easy Adaptation to New Hardware: We design our approach so that migrating to different input or output devices requires adjusting only a small set of physically interpretable parameters, ensuring robust adaptation across diverse hardware setups.
@@ II Related Work
Figure 3: Pipeline of Our Modules: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides visual preview before physical execution. We achieve precise alignment between virtual and physical robots through AprilTag calibration.
@@ II-A Robot Teleoperation Frameworks for Manipulation
The growing interest in training robots through imitation learning has driven the need for extensive, high-quality real robot datasets. Teleoperation provides an efficient method for demonstrating and recording intricate robotic tasks[13, 14, 15]. Researchers have explored various teleoperation systems, including VR controllers [1, 2, 3], motion capture [16, 17, 18], wearable gloves [19, 11, 20, 21], exoskeletons [22, 23], each offering distinct benefits in terms of accessibility, precision, and generalizability.
Vision-based teleoperation systems [6] typically employ RGB or RGB-D cameras to detect hand poses. The hand pose retargeting module maps the human hand pose data obtained from perception algorithms into joint positions of the dexterous hand. We formulate this process as an optimization problem that minimizes the difference between the key point vectors of the human hand and the dexterous hand. To generate the end-effector motions, 3D positions of key points in the camera frame are calculated by using an RGB-D camera. The wrist pose is then computed by keypoint positions in the local wrist frame and global camera frame using the Perspective-n-Point (PnP) algorithm.
Although vision-based teleoperation systems offer a relatively low-cost solution [6, 2, 3, 1], they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspace, and high computational demands. This limitation stems from the requirement of vision-based approaches for operators to keep their hands within the camera’s field of view, thereby restricting the action space of the human hand. Although increasing the scale factor can amplify end-effector movement and expand the action space of the robot, it aggravates the difficulty for operators in performing fine-grained tasks. These approaches also struggle to handle scenarios where fingers overlap. The extra keypoint retargeting module can be computationally expensive in real-time. Although specialized hardware [4] can improve accuracy, it substantially increases system costs. Our proposed system directly uses the joint pose obtained from the motion capture suit and data gloves, preventing the deviation of vision-based systems and reducing the computational cost. Recent unified teleoperation frameworks attempt to address these limitations but often introduce new trade-offs. Some approaches [3, 6] prioritize implementation simplicity by tracking only fingertip positions rather than enabling detailed joint-level control. Furthermore, your customization options typically require time-consuming manual parameter tuning that requires significant expertise.
@@ II-B Visual Feedback in Teleoperation Systems
Visual feedback plays a crucial role in teleoperation systems by providing users with spatial awareness and task-relevant information. Traditional approaches primarily rely on real-time camera feeds [24, 25], which can be limited by occlusions and restricted viewing angles. Moreover, the lack of preview capabilities for intended commands leaves users (especially new users) uncertain about the robot’s response, which can increase error rates and user fatigue.
@@
To address these limitations, researchers have explored enhanced visualization techniques [26, 27, 28, 29, 30] that overlay virtual information on the workspace view. Even though these approaches improve user perception by displaying planned trajectories and environmental constraints, they primarily design and visualize AR objects for a specific task, such as a pink box with dimensions of 10 cm × 10 cm × 10 cm. Any changes to the object’s attributes (e.g., dimensions, color, or type) invalidate the preconfigured AR object, so the system must rebuild it to match the new characteristics.
Instead of visualizing the manipulated objects, our system provides an AR representation of the robot that directly maps operator commands. This design ensures safety for new users who may be unfamiliar with how the robot responds to their inputs. Our configuration allows users to use their own URDF file, eliminating the need to reconstruct manipulated objects for every new task.
III System Overview
Our system offers an integrated workflow for controlling robot movements with both precision and safety, leveraging low-cost hardware (under \$1,000) and diverse input devices. This workflow comprises two complementary modules:
  1. 1.
    Teleoperation Module: Includes a teleoperation pipeline for robust motion retargeting.
  2. 2.
    Preview-Based Module: Allows operators to check intended movements in a virtual setting before execution.
As illustrated in Fig. 3, the teleoperation module (Section V) offers a safe and intuitive pipeline that proceeds in three stages:
  1. 1.
    Wrist pose estimation via IMU-based tracking (Section IV-A),
  2. 2.
    Hand pose estimation using a motion capture glove (Section IV-B),
  3. 3.
    Non-collision retargeting to ensure safe joint configurations (Section IV-C).
Together, these components collectively enable an efficient mapping from human motion to robot commands. Moreover, by adopting SMPL-X (Section IV-A1) as a standardized representation of the human body, our pipeline can seamlessly integrate data from any input device conforming to this format.
Additionally, as illustrated in Fig. 3, the preview-based module (Section V) then provides a safe and intuitive pipeline that proceeds in three stages:
  1. 1.
    The preview robot is aligned with the physical robot using AprilTag markers [31] (Section V-B1).
  2. 2.
    User commands are visualized on the aligned preview robot (Section V-B2).
  3. 3.
    Once the I/O state transitions from active to inactive, the final preview pose is captured and converted into an optimal trajectory for actual execution (Section V-B3).
This preview-then-execute scheme reduces errors and protects the robot and its surroundings by allowing users to visualize and refine their actions beforehand. Thus, it substantially improves efficiency and safety.
IV Teleoperation Pipeline
IV-A Wrist Pose Estimation
Our system tracks the human operator’s wrist motion to enable intuitive robot teleoperation. The input consists of raw IMU sensor data from sensors attached to the operator’s arm, while the output is the 6-DoF wrist pose (position and orientation) in world coordinates. This wrist pose serves as the primary control signal for the robot’s end-effector, allowing natural mapping between human arm movements and robot motion.
IV-A1 SMPL-X Standard
@@
We adopt the SMPL-X standard [32] as our standard representation of the human body. The SMPL-X model represents a kinematic tree with standardized joint coordinate systems, where each joint’s pose describes a rotation and translation relative to its parent frame. This hierarchical structure enables consistent pose representation across different motion capture devices and simplifies the integration with our TelePreview system.
IV-A2 IMU-based Tracking System
@@
IMU-based motion capture offers robust tracking regardless of visual conditions and occlusions. Following the SMPL-X convention, we define the world frame system at the midpoint between the feet. All positions and orientations in this paper are expressed in this coordinate system unless otherwise specified. After obtaining the SMPL-X model, we compute the wrist position $\mathbf{p}_{w}\in\mathbb{R}^{3}$ and orientation $\mathbf{R}_{w}\in SO(3)$ relative to this world frame.
@@
IV-B Hand Pose Estimation
Mocap gloves [33] offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors. Our joint-to-joint mapping approach provides a generalizable solution for retargeting high-dimensional human hand motion data to lower-dimensional robotic systems. This approach is particularly effective when the input space (e.g., the human hand with 27 DoF) has higher dimensionality than the target space (e.g., the robotic hand with 16 DoF), as it allows selective mapping of the most intuitive and task-relevant degrees of freedom. Through manual selection of corresponding joints based on human intuition and task requirements, followed by range alignment between input and output spaces, we maintain natural and predictable motion correspondence.
@@ IV-B Hand Pose Estimation
(a) Current Retargeting Method [6] Causes Collision.
@@
IV-C Non-Collision Retargeting Network
Direct mapping from human hand postures to robotic hands using existing retargeting approaches like [6] can result in self-collisions, as demonstrated in Fig. 4(a). While traditional retargeting methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations.
To address these challenges, we develop a learning-based framework with two key components:
  1. 1.
    A Self-Collision Prediction Network (CPN) that takes robot joint configurations $q\in\mathbb{R}^{n}$ as input and outputs a binary collision probability vector $p\in\mathbb{R}^{m}$ indicating collision likelihood for each link.
  2. 2.
    A Configuration Correction Network (CCN) that maps collision-prone joint configurations to their nearest collision-free counterparts. Specifically, it transforms an invalid robot configuration $q_{invalid}\in\mathbb{R}^{n}$ to a valid configuration $q_{valid}\in\mathbb{R}^{n}$ that minimizes both collision risk and deviation from the original pose.
As demonstrated in Fig. 4(b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz).
V Preview Pipeline
V-A Why Preview System Assistance?
@@
Traditional teleoperation approaches face significant challenges in data quality when collecting demonstrations for imitation learning. During complex manipulation tasks, users frequently make exploratory movements that contaminate the demonstration data with sub-optimal trajectories. These repeated probing actions to locate the socket edge and refine the approach angle before successful insertion introduced significant noise into the collected datasets. Such noise is particularly problematic for training imitation learning models, which benefit most from clean, purposeful demonstrations rather than trajectories cluttered with exploratory movements.
To address these limitations and provide a more efficient workflow, we propose incorporating a dedicated preview feature that allows users to refine their actions before physical execution. By separating exploration from the final command, our approach reduces sub-optimal motions that would otherwise degrade data quality.
V-B Technical Implementation of Preview System
We design our preview feature to integrate seamlessly with the teleoperation workflow. It requires three critical elements: (1) precise spatial calibration to maintain alignment between the virtual preview and the physical robot, (2) realistic rendering to blend the preview with camera feeds, and (3) state management to handle transitions between preview and execution phases. The following subsections detail each of these technical components, highlighting how they collectively enable users to view and adjust planned motions prior to committing them to the physical robot.
V-B1 Spatial Alignment Using AprilTag
To maintain a precise spatial alignment between the preview and the physical robot in the camera viewpoints, we implement a robust calibration system using AprilTags [31]. Our approach involves two sequential steps: (1) a hand-eye calibration that establishes the transformation between the robot’s base frame and the fixed camera frame, and (2) AprilTag pose detection that enables real-time tracking of the robot’s position and orientation relative to each camera view. This chain of transformations ensures consistent preview visualization across different viewpoints.